### 31 Jan 2024 v1

@@ 4 Discussion
The high and polarisation fraction seen in the strong activity period suggest a coherent emission mechanism at play, either via an ECME mechanism or a plasma emission process (see, Melrose, 1993, for an overview). The following subsections will summarise our current understanding of the two mechanisms.
@@ 4.1 Plasma emission
Plasma emission mechanism is the other coherent emission process which can also generate highly polarised emission (Tsytovich & Kaplan, 1969; Melrose & Sy, 1972). However, this mechanism is often subject to near-source scattering (e.g., Steinberg et al., 1971; Robinson, 1983; Arzner & Magun, 1999; Mohan et al., 2019; Kontar et al., 2019; Chen et al., 2020), which reduces the observed polarisation levels. In the case of solar type-III bursts, which are well known plasma emission sources, the maximum Stokes V/I detected is 60% (Saint-Hilaire et al., 2013; Rahman et al., 2020). However, these results are at frequencies below 100 MHz. Mohan (2021) showed that the effects of scattering could be less significant at higher frequencies since the effective scattering screen widths decreases towards higher frequency or equivalently at lower heights above corona. This means that the scattering induced depolarisation at our observing frequencies could be lower than that observed close to 100 MHz, but its effects may not be negligible. It is therefore unlikely that we observe radio bursts with polarisation levels above 75% generated purely via a coherent plasma emission mechanism.
4.2 Electron Cyclotron Maser Emission (ECME)
@@
An ECME is driven by a loss-cone distribution of electrons. Such an unstable electron pitch-angle distribution can form in post-flare coronal loops (e.g. Wu & Lee, 1979; Winglee & Dulk, 1986; Ning et al., 2021). The ECME emission mechanism predicts that in /¡0.3 X-mode (polarised in the sense of electron gyration) emission is dominant at fundamental . Recent studies using fully kinetic particle-in-cell simulations suggest that a harmonic X-mode emission via Z-mode coalescence can also form at /¡0.3 and be observable (Ning et al., 2021). Meanwhile, in the 0.3¡/¡1.3 range, Z mode emission at dominates (Melrose et al., 1984b; Winglee, 1985) with polarisation in the opposite sense of electron gyration (O-mode). But, X-mode emission at higher harmonics of can also form by the coalescence of Z-mode waves when 0.3¡/¡1.3 regime (Melrose, 1991). In a narrow window of 1.3¡/¡1.5 X-mode emission emitted directly at the harmonics of can be observed (Melrose et al., 1984a). X-mode emission at harmonics of are expected via either a direct harmonic emission or various coalescence interactions up to /¡3 (Bastian et al., 1990). Harmonics higher than three are hardly produced via any of these process due to the requirement of very high energy density in the unstable wave modes (Melrose, 1991). So 3 is the highest observable harmonic expected. Another crucial factor is also that the pure X-mode emission at , or an X-mode emission at harmonics of , formed by Z-mode coalescence can be nearly 100% polarised (Melrose, 2009, 2017b).
@@
When /¿ 3, the emissivity of the various ECME processes drops dramatically and along with the attenuation due to gyromagnetic absorption, ECME is expected to be insignificant (Melrose & Dulk, 1982; Bastian, 1990; Melrose, 1991; Ning et al., 2021).
@@
To summarise, the possible X-mode polarised ECME pathways driven by loss-cone electron distribution are:
@@
0.3¡/¡1.3: Z/W wave coalescence can produce X-mode emission at s.
@@
Beyond /=3, plasma emission mechanism dominates. Ni et al. (2020) showed that when /, the Z, W and UH modes triggered by an electron cyclotron maser instability would not escape the plasma. Instead these waves can interact via a plasma emission mechanism to generate coherent emission at the fundamental and harmonic of the upper hybrid frequency, .
Besides, the existing theory suggests that O-mode emission can be produced via the following pathway:
1.
0.3¡/¡1.3: Z-mode pathway generates emission at .
@@ 4.3 The emission mechanism of AD Leo flares
In this subsection, we will discuss the various possible mechanisms for the observed emission given the possible local physical scenarios, during the various activity periods. First, we build a 1D density and magnetic field model for the star, using which and will be estimated across coronal heights. Zeeman Doppler Imaging (ZDI) based studies (Morin et al., 2008; Bellotti et al., 2023) infer that the magnetic field of AD Leo is predominantly dipolar () with surface field strengths within 0.1 - 0.5 kG. However, at strongly flaring active regions magnetic field strengths on AD Leo can be as high as 3kG (e.g., Cranmer & Saar, 2011). Based on this, we choose a typical range for field strength at the base of the active region, , between 0.3 and 1 kG. The chosen range considers a mean ZDI field strength at the lower end and the order of magnitude of the strongest reported active region —B— at the higher end. For the 1D analytical model of the radial magnetic field, two models are considered. First one is a dipole-dominated active region magnetic field model suggested for solar active regions by Aschwanden et al. (1999),
@@
with the dipole scale height, assumed to be 0.1R (R is the stellar radius), based on solar studies (Aschwanden et al., 1999). Note that r is measured radially from stellar centre in units of R. In the second model, we assume a simple multipole expansion model of magnetic field with the net magnetic flux assumed to be distributed as 72% dipole, 21% quadrupole and 7% octapole fraction, based on the ZDI results for observations done in 2020 (Bellotti et al., 2023).
@@
The fractional composition of the magnetic moments employed here, are based on large scale ZDI results that may not be representative of flaring active regions. Since there exist no estimation of multipole moments at the stellar active region, we assume the large-scale fractional distribution. However, the high relative dipolar concentration in Eqn. 2 assures that this is probably not an uncanny assumption for an active region magnetic structure if the AD Leo active regions are also similar to those on the sun with strong bipolar magnetic field concentrations and arcade-like structures. The two models will be referred to as Model 1 (Eqn. 1) and Model 2 (Eqn. 2) hereafter.
We derive assuming the density model () proposed by Villadsen & Hallinan (2019),
@@
The basal coronal density in the model is derived from OVII line ratios by Ness et al. (2004). X-ray observations of pre-flare loops also suggest a basal coronal density within 10 - 10 (Namekata et al., 2020). Also, the assumed radial density profile is empirical, but consistent with the best-fit hydrostatic equilibrium profiles derived for solar active regions, based on direct measurements (e.g. Mann et al., 1999; Aschwanden et al., 2001). Zucca et al. (2014) showed that the data from a CME driving active solar region is well modelled by an exponential hydrostatic equilibrium density profile for r, which is the height range of interest in our study.
@@
Figure 4(a) shows the derived and . The plotted range of correspond to 0.31 kG. The observing band (550 - 850 MHz) is marked in red. The green band shows the frequencies corresponding to one third of the observation band. First, consider Model 1 for the observing band. For an ECME emission process, the observing frequency, = s where s =1,2,3. The region where the red band overlaps the blue corresponds to the r range where =. Similarly, the r range where the green band overlaps the blue identifies the r range where =3. Beyond this r range, the possibility of ECME to be observed in 500 - 850 MHz is very low since inherent emissivity drops significantly (Melrose et al., 1984b). Vertical solid red lines mark the extent of the overlapping r range for =, and the solid green lines mark the r range for =3. Same exercise is performed for Model 2. Instead of solid lines, we use dotted lines of the same colour to mark the r ranges.
@@
Figure 4(b) shows the ratio / for = , and /3. Horizontal lines mark /=0.3 ,1.3 and 3. We will first consider Model 1. The overlap between the red and blue bands give us the / values for the r range from where fundamental ECME mode emission can originate in the observing band. Clearly this ratio is greater than 1.3 but less than 3, implying that the fundamental X-mode emission is impossible but a Z-mode ECME is possible. However, the polarisation will be in the opposite sense of electron gyration (O-mode). However, an X-mode polarised emission is possible via the coalescence of Z or W-waves, as per the standard theory (Melrose, 1991). But, the resultant harmonic X-mode emission would not be observable in our band, but in 2 band. Meanwhile, the overlap between green and blue bands give us the / values in the r range from where if third harmonic emission arises, it can be observed in our . The range of / is much greater than 3, making ECME emission mechanism unfavourable. Similar to the above discussion, for Model 2, the intersection between gray and red band gives the range of / for fundamental ECME emission and the overlap between gray and green band gives the range of the ratio for harmonic emission to be observable in band. Consider the gray and red band overlap. We have cases for 0.3¡/¡3. Either a O or Z mode emission (0.3¡/¡1.3) or a harmonic X-mode (1.3¡/¡3) can be triggered here. But, akin to the case of Model 1, the harmonic X-mode emission born out of Z/W-mode coalescence or a direct emission cannot be observed in our observing band. However, the overlap between green (= /3) and gray band falls within 1.3¡/¡3 favouring a harmonic X-mode ECME from its r range via Z or W-mode coalescence or a direct harmonic emission, observable in band.
To summarise the discussion:
1.
Model 1 only favours an O-mode polarised (opposite sense of electron gyration) emission at via Z-mode pathway, observable in the band
2.
@@
Model 2 favours an X-mode (polarised in the sense of electron gyration) emission at harmonics of local formed via either a direct emission or Z/W-mode coalescence pathways, and observable in band.
@@
We now consider the polarisation in the bursts to further constrain the mechanism. The polarisation has consistently been left circular with Stokes V/I ¿ 60% during bursts. At higher frequencies (700 MHz), we find higher polarisation degrees attained during strong flares (e.g., Fig. 3). Morin et al. (2008) had shown that the south pole of AD Leo is inclined well along the line of sight (LOS) at an angle of 20. So there is a high chance that in the LOS magnetic field structures could be oriented southwards. This means the observed emission is in the sense of electron gyration,which means the emission is in X-mode. Also, the degree of polarisation is during flares, often reaching 100% at higher frequencies (see Fig. 3). If the emission is via an ECME process, the possible pathways are either a fundamental X-mode ECME mechanism from regions where = or a harmonic X-mode emission from regions where =/s and 0.3¡/¡3 (s=2,3). Based on the inferences from Fig. 4, Model 1 does not favour both fundamental or harmonic X-mode ECME, but Model 2 favours a harmonic X-mode emission by all possible pathways. Meanwhile, Zhang et al. (2023) observed the star in the 1 - 1.5 GHz band during the strong activity period, between 21:13 – 22:48 UT, and reported emission features that are better explained by an ECME mechanism. Unfortunately their observations did not extend further into the weak and quiescent activity phases.
@@
However, this analysis does not explicitly consider the propagation effects that would affect the observed ECME. Without multi-waveband spectroscopic data, the existing stellar atmospheric models (e.g. Fontenla et al., 2016; Wedemeyer et al., 2013) cannot be well constrained to derive a reliable ambient/quiescent physical structure (especially, B(r) and ), crucial to assess the atmospheric opacity (see White & Kundu, 1997; Treumann, 2006, for an overview). Consequently, based on current knowledge on the ECME mechanism as derived fromsolar observations, the emissivity should fall drastically above (see, Sect. 4.2). Moreover, the emission for higher harmonics (s=) is less affected by gyromagnetic absorption from upper coronal layers (Zheleznyakov, 1970; Melrose & Dulk, 1982). Both these strongly favour the detection of ECME at s=2,3. Our inferences on ECME generation based on 1D stellar models also favour harmonic emission.
@@
Plasma emission, which is the competing mechanism to ECME produces coherent emission observable at (F-branch) and 2 (H-branch) (Ginzburg & Zhelezniakov, 1958; Tsytovich & Kaplan, 1969; Melrose & Sy, 1972; Melrose, 2009). This mechanism generally favours polarisation in the O-mode (Melrose, 2017b) especially for the F branch. For the H branch, this is true if , but if /1.5, highly polarised X-mode emission at =2 is possible (Willes & Melrose, 1997). Given that we observe a high X-mode polarised flares, it is likely a H-branch emission from regions with /1.5. From Fig. 5b, we infer that the H-band emission observable in should arise from r1.9R, where ¡/2. In this r range, ¿¿ for Model 1. But for some B values, /¡1.5 is satisfied for Model 2. So only Model 2 can support a H-band plasma emission in X-mode.
@@
Based on the discussions so far, Model 2 favours X-mode polarised emission via both H-branch plasma emission mechanism and ECME emission processes at harmonics of . Note that Model 1 does not support any of the likely emission processes favouring a strong X-mode ECME or H-band plasma emission. This possibly means that a solar like active region model is not good for M dwarfs, but a ZDI based multipole expansion model works better. For the rest of the discussion, we will hence consider only Model 2. Meanwhile, the H-branch plasma emission observed in the sun has always shown polarisation levels below 30% (Willes & Melrose, 1997; Saint-Hilaire et al., 2013; Rahman et al., 2020). So, during strong activity period, given the high polarisation of 60 - 100% (Figs. 1-3) the emission mechanism is likely via an ECME mechanism.
@@
In the analysis so far, we had kept the basal coronal density constant in the model. Let us consider a variation of this value within an order of magnitude and assess the possible ECME pathways supported by the two B(r) models. For this analysis, we vary n within 10 - 10 cm, consistent with the values reported by previous spectroscopic and X-ray continuum observations. This corresponds to a variation in the assumed basal coronal density in Eq. (3) by 0.4 to 4 times. If the basal density goes up 4 times, since we would see a an increase in by a factor of 2, which in turn raises all the / bands in Fig. 4 by the same factor. Note that in Fig. 4 the green band overlaps with gray (Model 2) within 1¡/¡3. So the harmonic X-mode emission pathway, previously supported by Model 2 will remain feasible but now in the 2¡/¡3. The other ECME pathways that were already not supported by the previously assumed lower density plasma will anyway be unsupported by a denser plasma due to enhanced damping of the ECME driving instability. However, since /3 overlaps with Model 2 in /¡10, an ECME driven coherent plasma emission process suggested by Ni et al. (2020) can produce harmonic X-mode emission observable in . Meanwhile, if we assume a 0.4 times lower density in the active region, all the / bands in Fig. 4 will be reduced by a factor of 1.6. In this case for Model 2, harmonic X-mode emission will be observable as earlier. Interestingly, a portion of the overlapping region between the Model 1 (blue band) the green band will fall within 2.5¡/¡3, suggesting a possibility of observing a harmonic X-mode emission in . This may make Model 1 viable for a small range of B and basal density values. However, Model 2 consistently supports an ECME mechanism for an order of magnitude wide range of B and n values.
4.4 Weak activity period - type-IV burst

### 04 Mar 2024 v2

@@ 4 Discussion
The high and polarisation fraction seen in the strong activity period suggest a coherent emission mechanism is at play, either via an ECME mechanism or a plasma emission process (see, Melrose, 1993 for an overview). The following subsections summarise our current understanding of the two mechanisms.
@@ 4.1 Plasma emission
A plasma emission mechanism is the other coherent emission process, which can also generate highly polarised emission (Tsytovich & Kaplan, 1969; Melrose & Sy, 1972). However, this mechanism is often subject to near-source scattering (e.g. Steinberg et al., 1971; Robinson, 1983; Arzner & Magun, 1999; Mohan et al., 2019; Kontar et al., 2019; Chen et al., 2020), which reduces the observed polarisation levels. In the case of solar type-III bursts, which are well-known plasma-emission sources, the maximum Stokes V/I detected is 60% (Saint-Hilaire et al., 2013; Rahman et al., 2020). However, these results are at frequencies below 100 MHz. Mohan (2021) showed that the effects of scattering could be less significant at higher frequencies since the effective scattering screen widths decrease towards higher frequency or equivalently at lower heights above the corona. This means that the scattering-induced depolarisation at our observing frequencies could be lower than that observed close to 100 MHz, but its effects may not be negligible. It is therefore unlikely that we observe radio bursts with polarisation levels above 75% generated purely via a coherent plasma emission mechanism.
4.2 Electron cyclotron maser emission
@@
An ECME is driven by a loss-cone distribution of electrons. Such an unstable electron pitch-angle distribution can form in post-flare coronal loops (e.g. Wu & Lee, 1979; Winglee & Dulk, 1986; Ning et al., 2021). The ECME emission mechanism predicts that in /¡0.3 X-mode (polarised in the sense of electron gyration), emission is dominant at fundamental . Recent studies using fully kinetic particle-in-cell simulations suggest that a harmonic X-mode emission via Z-mode coalescence can also form at /¡0.3 and be observable (Ning et al., 2021). Meanwhile, in the 0.3¡/¡1.3 range, Z-mode emission at dominates (Melrose et al., 1984b; Winglee, 1985) with polarisation in the opposite sense of electron gyration (O-mode). However, X-mode emission at higher harmonics of can also form by the coalescence of Z-mode waves when there is a 0.3¡/¡1.3 regime (Melrose, 1991). In a narrow window of 1.3¡/¡1.5, X-mode emission emitted directly at the harmonics of can be observed (Melrose et al., 1984a). X-mode emission at harmonics of are expected via either a direct harmonic emission or various coalescence interactions up to /¡3 (Bastian et al., 1990). Harmonics higher than three are hardly produced via any of these process due to the requirement of very high energy density in the unstable wave modes (Melrose, 1991). So, 3 is the highest observable harmonic expected. Another crucial factor is that the pure X-mode emission at , or an X-mode emission at harmonics of , formed by Z-mode coalescence can be nearly 100% polarised (Melrose, 2009, 2017b).
@@
When /¿ 3, the emissivity of the various ECME processes drops dramatically, and along with the attenuation due to gyromagnetic absorption, ECME is expected to be insignificant (Melrose & Dulk, 1982; Bastian, 1990; Melrose, 1991; Ning et al., 2021).
@@
To summarise, the possible X-mode polarised ECME pathways driven by loss-cone electron distribution are as follows:
@@
0.3¡/¡1.3: Z or W wave coalescence can produce X-mode emission at s.
@@
Beyond /=3, a plasma emission mechanism dominates. Ni et al. (2020) showed that when /, the Z, W, and UH modes triggered by an electron cyclotron maser instability would not escape the plasma. Instead, these waves can interact via a plasma emission mechanism to generate coherent emission at the fundamental and harmonic of the upper hybrid frequency, .
Besides this, the existing theory suggests that O-mode emission can be produced via the following pathway: 0.3¡/¡1.3: Z-mode pathway generates emission at .
@@ 4.3 The emission mechanism of AD Leo flares
In this subsection, we discuss the various possible mechanisms for the observed emission given the possible local physical scenarios during the various activity periods. First, we built 1D radial models of density and magnetic field for the star and used them to estimate and across coronal heights. Zeeman-Doppler imaging (ZDI) based studies (Morin et al., 2008; Bellotti et al., 2023) infer that the magnetic field of AD Leo is predominantly dipolar (), with surface field strengths within 0.1 - 0.5 kG. However, the magnetic field strengths at strongly flaring active regions can be as high as 3kG (e.g. Cranmer & Saar, 2011). Based on this, we chose a typical range for field strength at the base of the active region, , between 0.3 and 1 kG. The chosen range considers a mean ZDI field strength at the lower end and the order of magnitude of the strongest reported active region —B— at the higher end. For the 1D analytical model of the radial magnetic field, two models are considered. The first one is a dipole-dominated active region magnetic field model suggested for solar active regions by Aschwanden et al. (1999):
@@
with the dipole scale height, assumed to be 0.1R (R is the stellar radius), based on solar studies (Aschwanden et al., 1999). We note that r is measured radially from stellar centre in units of R. In the second model, we assume a simple multipole expansion model of magnetic field with the net magnetic flux assumed to be distributed as 72% dipole, 21% quadrupole, and 7% octapole fraction, based on the ZDI results for observations from 2020 (Bellotti et al., 2023):
@@
The fractional composition of the magnetic moments employed here are based on large-scale ZDI results that may not be representative of flaring active regions. Since no estimation of multipole moments at the stellar active region exists, we assume the large-scale fractional distribution. However, the high relative dipolar concentration in Eq. 4.3 assures that this is probably not an uncanny assumption for an active region’s magnetic structure if the AD Leo active regions are also similar to those of the Sun with strong bipolar magnetic field concentrations and arcade-like structures. The two models are referred to as Model 1 (Eqn. 1) and Model 2 (Eqn. 4.3) hereafter.
We derive assuming the density model () proposed by Villadsen & Hallinan (2019):
@@
The basal coronal density in the model is derived from OVII line ratios by Ness et al. (2004). X-ray observations of pre-flare loops also suggest a basal coronal density within 10 - 10 (Namekata et al., 2020). Meanwhile, the assumed radial density profile is empirical, but consistent with the best-fit hydrostatic equilibrium profiles derived for solar active regions based on direct measurements (e.g. Mann et al., 1999; Aschwanden et al., 2001). Zucca et al. (2014) showed that the data from a solar region that hosted a CME, is modelled well by an exponential hydrostatic equilibrium density profile for r, which is the height range of interest in our study.
@@
Figure 4(a) shows the derived and . The plotted range of corresponds to 0.31 kG. The observing band (550 - 850 MHz) is marked in red. The green band shows the frequencies corresponding to one third of the observation band. For an ECME emission process, the observing frequency = s, where s =1,2,3. First, we considered Model 1 (blue band). The region where the red band overlaps the blue one corresponds to the r range where =. Similarly, the r range where the green band overlaps the blue one identifies the r range where =3. Beyond this r range, the possibility of observing ECME in 500 - 850 MHz is very low since inherent emissivity drops significantly (Melrose et al., 1984b). Vertical solid red lines mark the extent of the overlapping r range for =, and the solid green lines mark the r range for =3. The same exercise is performed for Model 2 (grey band). Instead of solid lines, we used dotted lines of the same colour to mark the r ranges.
@@
Figure 4(b) shows the ratio / for = , and /3. Horizontal lines mark /=0.3, 1.3, and 3. We first consider Model 1. The overlap between the red and blue bands gives us the / values for the r range from where fundamental ECME mode emission can originate in the observing band. Clearly this ratio is greater than 1.3, but less than 3, implying that the fundamental X-mode emission is impossible, but a Z-mode ECME is possible. However, the polarisation will be in the opposite sense of electron gyration (O-mode). Moreover, an X-mode polarised emission is possible via the coalescence of Z or W waves, as per the standard theory (Melrose, 1991). Yet, the resultant harmonic X-mode emission would not be observable in our band, but in the 2 band. Meanwhile, the overlap between green and blue bands gives us the / values in the r range from where, if third harmonic emission arises, it can be observed in our . The range of / is much greater than three, making an ECME emission mechanism unfavourable. Similar to the above discussion, for Model 2, the intersection between the grey and red bands gives the range of / for fundamental ECME emission, and the overlap between the grey and green bands gives the range of the ratio for harmonic emission to be observable in band. If we consider the grey and red band overlap, we have cases for 0.3¡/¡3. Either an O or Z mode emission (0.3¡/¡1.3) or a harmonic X-mode (1.3¡/¡3) can be triggered here. However, akin to the case of Model 1, the harmonic X-mode emission born out of Z- or W-mode coalescence or a direct emission cannot be observed in our observing band. However, the overlap between the green (= /3) and grey bands falls within 1.3¡/¡3, favouring a harmonic X-mode ECME from its r range via Z- or W-mode coalescence or a direct harmonic emission, which is observable in the band.
@@
To summarise , Model 1 only favours an O-mode polarised (opposite sense of electron gyration) emission at via a Z-mode pathway, which is observable in the band. However, Model 2 favours an X-mode (polarised in the sense of electron gyration) emission at harmonics of local formed via either a direct emission or Z/W-mode coalescence pathways and observable in the band.
@@
We now consider the polarisation in the bursts to further constrain the mechanism. The polarisation has consistently been left circular with Stokes V/I ¿ 60% during bursts. At higher frequencies (700 MHz), we find higher polarisation degrees attained during strong flares (e.g. Fig. 3). Morin et al. (2008) had shown that the south pole of AD Leo is inclined well along the line of sight (LOS) at an angle of 20. So, there is a high chance that magnetic field structures could be oriented southwards on the LOS. This means the observed emission is in the direction of electron gyration,which means the emission is in X-mode. Also, the degree of polarisation is during flares, often reaching 100% at higher frequencies (see Fig. 3). If the emission is via an ECME process, the possible pathways are either a fundamental X-mode ECME mechanism from regions where = or a harmonic X-mode emission from regions where =/s and 0.3¡/¡3 (s=2,3). Based on the inferences from Fig. 4, Model 1 does not favour both fundamental and harmonic X-mode ECME, but Model 2 favours a harmonic X-mode emission by all possible pathways. Meanwhile, Zhang et al. (2023) observed the star in the 1 - 1.5 GHz band during the strong activity period, from 21:13 – 22:48 UT, and reported emission features that are better explained by an ECME mechanism. Unfortunately, their observations did not extend further into the weak and quiescent activity phases.
@@
Our analysis did not explicitly consider the propagation effects that would have affected the observed ECME. Without multi-waveband spectroscopic data, the existing stellar atmospheric models (e.g. Fontenla et al., 2016; Wedemeyer et al., 2013) cannot be well constrained to derive a reliable ambient/quiescent physical structure (especially B(r) and ), which is crucial to assessing the atmospheric opacity (see White & Kundu, 1997; Treumann, 2006). Consequently, based on current knowledge on the ECME mechanism as derived from solar observations, the emissivity should fall drastically above (see Sect. 4.2). Moreover, the emission for higher harmonics (s=) is less affected by gyromagnetic absorption from upper coronal layers (Zheleznyakov, 1970; Melrose & Dulk, 1982). Both of these strongly favour the detection of ECME at s=2,3. Our inferences on ECME generation based on 1D stellar models also favour harmonic emission.
@@
Plasma emission, which is the competing mechanism to ECME, produces coherent emission observable at (F-branch) and 2 (H-branch) (Ginzburg & Zhelezniakov, 1958; Tsytovich & Kaplan, 1969; Melrose & Sy, 1972; Melrose, 2009). This mechanism generally favours polarisation in the O-mode (Melrose, 2017b) especially for the F branch. For the H branch, this is true if , but if /1.5, highly polarised X-mode emission at =2 is possible (Willes & Melrose, 1997). Given that we observe a high X-mode polarised flares, it is likely a H-branch emission from regions with /1.5. From Fig. 5b, we infer that the H-band emission observable in should arise from r1.9R, where ¡/2. In this r range, ¿¿ for Model 1. However, for some B values, /¡1.5 is satisfied for Model 2. So, only Model 2 can support an H-band plasma emission in X mode.
@@
Based on the discussions so far, Model 2 favours X-mode polarised emission via both H-branch plasma emission mechanisms and ECME emission processes at harmonics of . We note that Model 1 does not support any of the likely emission processes favouring a strong X-mode ECME or H-band plasma emission. This possibly means that a solar-like active region model is not good for M dwarfs, but a ZDI-based multipole expansion model works better. For the rest of the discussion, we hence only consider Model 2. Meanwhile, the H-branch plasma emission observed in the Sun has always shown polarisation levels below 30% (Willes & Melrose, 1997; Saint-Hilaire et al., 2013; Rahman et al., 2020). So, during a strong activity period, given the high polarisation of 60 - 100% (Figs. 1-3), the emission mechanism is likely via an ECME mechanism.
@@
In the analysis so far, we had kept the basal coronal density constant in the model. We now consider a variation of this value within an order of magnitude and assess the possible ECME pathways supported by the two B(r) models. For this analysis, we vary n within 10 - 10 cm, consistent with the values reported by previous spectroscopic and X-ray continuum observations. This corresponds to a variation in the assumed basal coronal density in Eq. (3) by 0.4 to 4 times. If the basal density goes up four times, since we would see an increase in by a factor of two, which in turn raises all the / bands in Fig. 4 by the same factor. We note that in Fig. 4 the green band overlaps with the grey one (Model 2) within 1¡/¡3. So, the harmonic X-mode ECME emission pathway, previously supported by Model 2, will remain feasible, but now in the 2¡/¡3. The other ECME pathways that were already not supported by the previously assumed lower density plasma will be unsupported by a denser plasma due to enhanced damping of the ECME driving instability. However, since /3 overlaps with Model 2 in /¡10, an ECME-driven coherent plasma emission process suggested by Ni et al. (2020) can produce harmonic X-mode emission observable in . Meanwhile, if we assume a 0.4 times lower density in the active region, all the / bands in Fig. 4 will be reduced by a factor of 1.6. In this case for Model 2, harmonic X-mode emission will be observable as before. Interestingly, a portion of the overlapping region between Model 1 (blue band) and the green band will fall within 2.5¡/¡3, suggesting a possibility of observing a harmonic X-mode emission in . This may make Model 1 viable for a small range of B and basal density values. However, Model 2 consistently supports an ECME mechanism for an order of magnitude range of B and n values.
4.4 Weak activity period and the type-IV burst